In the amusement park ride known as Magic Mountain Superman, powerful magnets accelerate a car and its riders from rest to 45.7 m/s in a time of 7.17 s. The mass of the car and riders is 6.15 × 103 kg. Find the average net force exerted on the car and riders by the magnets.
Answer #1

vf = vi + at

45.7 m/s = 0 + a(7.17 s)

a = 6.37 m/s2

Favg = ma

= (6.15 × 103 kg)(6.37 m/s2)

= 39198.74 N

***********************************************************************************************
This concludes the answers. If there is any mistake or omission, let me know immediately and I will fix it....